OpenStudy (anonymous):

If m<1=7x and m<2=5x+30 lines m and n are parallel when x equals 1) 12.5 2) 87.5 3) 15 4) 105 (explain)

OpenStudy (anonymous):



OpenStudy (nincompoop):

is that a less than or angle?

OpenStudy (nincompoop):

m and n are parallel if 1 and 2 are equal

OpenStudy (raden):

m<1 and m<2 forms alternate interior angles. they are same

OpenStudy (nincompoop):

that is cute to know, but solving for x is what proves angle 1 and angle 2 are the same

OpenStudy (nincompoop):

we are given y = 7x y = 5x + 30 from the looks of the equation, our slopes are not equal. well we can easily spot that, but we have to make it equal and that is the point of this exercise. so, we do the following: -7x + y = 0 (eq 1) -5x + y = 30 (eq 2) if we multiply eq 1 by -1 y becomes negative and then x becomes positive 7x - y = 0 -5x + y = 30 -------------- 2x = 30 x = 15

OpenStudy (nincompoop):

(eq 1) y = 7*15 = 105 (eq 2) y = 5*15 + 30 = 105 105 = 105